OpenStudy (anonymous):

determine if the series ((n^2+1)/(2n^2+1))^2n is absolutely convergent determine if the series ((n^2+1)/(2n^2+1))^2n is absolutely convergent @Mathematics

OpenStudy (anonymous):

\[((n^2+1)/(2n^2+1))^{2n}\]

OpenStudy (zarkon):

this \[\sum_{n=1}^{\infty}\left(\frac{n^2+1}{2n^2+1}\right)^{2n}\]

OpenStudy (zarkon):

if so use the root test

myininaya (myininaya):

\[\sum_{}^{}a_n \] \[\lim_{n \rightarrow \infty}|a_n|^\frac{1}{n} \] if less than one, then abs converges if greater than one, then diverges if =1, then inconclusive ( a waste of time and we are all very sad)

myininaya (myininaya):

so whats the limit of the inside if n goes to infinity?

myininaya (myininaya):

since the degrees inside are the same on both top and bottom then the limit inside is ( ? )^2 ^ | put your answer to my question here